OpenStudy (anonymous):

What is the base of a triangle with height = 7.4 cm and area = 48.1 cm2?

OpenStudy (anonymous):

A = 1/2 bh 48.1 = 1/2 (b)(7.4) get b alone 48.1 (2) = b(7.4) 96.2 = b(7.4) 96.2/7.4 = b b = 13

OpenStudy (anonymous):

area=1/2*base*height b=48.1(2)/7.4=13